OpenStudy (anonymous):

What is the simplified form of the expression? (3x^7/2)^6(x^2)^6

OpenStudy (chaise):

I'm assuming you are multiplying the two together, remember: \[a^m \times a^n=a ^{m+n}\]

OpenStudy (anonymous):

okay so i did it and my answer that i got is 3x^29. Is that right ?

OpenStudy (unklerhaukus):

\[\left(\frac{3x^7}{2}\right)^6(x^2)^6\] ?

OpenStudy (unklerhaukus):

\[a^m\times b^m=(a\times b)^m\]

OpenStudy (anonymous):

729x^29?

OpenStudy (unklerhaukus):

\[\left(3x^{7/2}\right)^6(x^2)^6\] ?